Using the information provided below, construct a linear model to maximize the profits for Puck a...

  1. Using the information provided below, construct a linear model to maximize the profits for Puck and Pawn Company that produces hockey sticks and chess sets.

Product

Labor

(Hr./Unit)

Wood

(Lb./Unit)

Profit

($/Unit)

Hockey Sticks (x1)

1

4

40

Chess Sets (x2)

2

3

50

Resource Availability:   40 hrs of labor per day; 120 lbs of wood

Decision Variables:       x1 = number of Hockey Sticks to produce; x2 = number of Chess Sets to produce

Identify the Objective Function:    Z = ___40____x1 + __50_____x2

Identify the Constraint Equation for Labor:    ___1____x1 + __2_____x2 <= ____40___

Identify the Constraint Equation for Wood:    _____4__x1 + ____3___x2 <= 120_______

Identify any other Constraint Equation(s):                           

Point A: X1 = ___30______ X2 = ______0___ Z = ____1200_____

Point B: X1 = _____24____ X2 = ____8_____ Z = _____1360____

Point C: X1 = ______0___ X2 = ___20______ Z = _____1000____

A.) Optimal point is (Circle one)   Point A       Point B       Point C

B.) Is there any slack associated with labor? Yes No, If yes, how many hours? _________

C.) Is there any slack associated with wood? Yes No, If yes, how many pounds? _________
Answer #1

Objective Function: Z = 40x1 + 50x2

Constraints:

x1 + 2x2 <= 40 [Constraint on Wood]

4x1 + 3x2 <= 120 [Constraint on labor]

x1 >= 0 [number of hockey sticks must be greater than zero]

x2 >= 0 [number of chess sets must be greater than zero]

We need to maximize objective function Z with the above constraints

a) Optimal Point will be the point B producing maximum profit of $1360

b) Number of labor hour used = 24 * 1 + 2 * 8 = 40

Available hours = 40

Hence there is no labor hours left, so the slack is (40-40) which is zero

No slack

c) Number of wood used = 24 * 4 + 3 * 8 = 120

Available hours = 120

Hence there is no wood hours left, so the slack is (120-120) which is zero

No slack